# Practical hybrid decoding scheme for parity-encoded spin systems

### v3

@@ V.1 Relationship between MCMC-BF hybrid algorithm and other known BF algorithms
Both the MCMC and BF decoding algorithms are hard-decision algorithms, in which only spin variables are of concern. The MCMC-BF decoding algorithm is considered an extended BF algorithm that incorporates MCMC decoding as a preprocessing step. We discuss the limitations of BF decoding and how they were resolved in the existing extended BF algorithms.
@@
Many researchers have developed sophisticated versions of the weighted BF (WBF) to incorporate reliability information of the hard decisions into the BF algorithm. See details in Refs.[36, 37] , and references therein. Alternatively, there is another approach to incorporate soft information. For example, Wadayama et al. proposed the Gradient Descent Bit Flipping (GDBF) algorithm [28]. They are based on modifications to the inversion function that may improve decoding performance, though this comes at the cost of increased decoding complexity. For example, the inversion functions employed in the BF, WBF, and GDBF algorithms are formally written as
$\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})=1+\sum_{i\in M(k)}s_{i}(\boldsymbol{x}),$ (52)
$\varDelta_{k}^{\mathrm{(WBF)}}(\boldsymbol{x})=\beta|J_{k}|+\sum_{i\in M(k)}w_{k}s_{i}(\boldsymbol{x}),$ (53)
@@
$\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})=J_{k}x_{k}+\sum_{i\in M(k)}s_{i}(\boldsymbol{x}),$ (54)
@@
respectively. Here, $\boldsymbol{x}=(x_{1},\ldots,x_{N_{v}})\in\{\pm 1\}^{N_{v}}$ is the current decision for the spin variables (VNs), $\beta$ is a positive real parameter to be adjusted,
$s_{i}(\boldsymbol{x})=\prod_{j\in N(i)}x_{j}\in\{\pm 1\}$ (55)
@@
is an $i$th syndrome (CN) for the decision $\boldsymbol{x}$ in the spin representation, and $J_{k}$ is identified with channel observation $y_{k}$ in the AWGN channel model. Here, the set $N(i)=\left\{j:H_{ij}=1\right\}$ is the VNs adjacent to an $i$th CN $\left(1\leq i\leq N_{c}\right)$ and the set $M(j)=\left\{i:H_{ij}=1\right\}$ is the CNs adjacent to a $j$th VN $\left(1\leq j\leq N_{v}\right)$. Please refer to Fig.2 for the definition of these sets. Note that as long as $s_{i}\left(\boldsymbol{x}\right)$ is a weight-3 syndrome, $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ is symmetric for the permutation of the elements of $\boldsymbol{x}$. In other words, $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ is invariant under exchanging $i\longleftrightarrow j$ for any pair of indices of VN other than $k$, justifying the all-one code-state assumption. In contrast, $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ is not symmetric if $s_{i}\left(\boldsymbol{x}\right)$ is a weight-4 syndrome. Similarly, because $\varDelta_{k}^{\mathrm{(WBF)}}(\boldsymbol{x})$ involves weight $w_{k}$ in the second term and $\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})$ involves $J_{k}$ in the first term, both depend on $\boldsymbol{J}$. Thus, $\varDelta_{k}^{\mathrm{(WBF)}}(\boldsymbol{x})$ and $\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})$ are not symmetric even if $s_{i}(\boldsymbol{x})$ is a weight-3 syndrome. It is interesting to note that the inversion functions in Eq.(52)-(54) can be formally derived from the following Hamiltonians:
$H^{\mathrm{(BF)}}(\boldsymbol{x})=-\sum_{i=1}^{N_{v}}x_{i}-\sum_{i=1}^{N_{c}}s_{i}(\boldsymbol{x}),$ (56)
$H^{\mathrm{(WBF)}}(\boldsymbol{x})=-\beta\sum_{i=1}^{N_{v}}\left|J_{i}\right|x_{i}-\sum_{i=1}^{N_{c}}w_{i}s_{i}(\boldsymbol{x}),$ (57)
$H^{\mathrm{(GDBF)}}(\boldsymbol{x})=-\frac{1}{2}\sum_{i=1}^{N_{v}}J_{i}x_{i}-\sum_{i=1}^{N_{c}}s_{i}(\boldsymbol{x}),$ (58)
@@
respectively. If we note that when we invert the sign of $x_{k}$, that is, $x_{k}\rightarrow-x_{k}$, the increase in energy $\Delta H_{k}^{\mathrm{(X)}}(\boldsymbol{x})$ is given by
$\Delta H_{k}^{(X)}(\boldsymbol{x})=2\varDelta_{k}^{(X)}(\boldsymbol{x}),$ (59)
@@
where $X=\mathrm{BF}$, $\mathrm{WBF}$, or $\mathrm{GDBF}$. It follows that if $\varDelta_{k}^{(X)}(\boldsymbol{x})<0$, flipping the $k$th spin reduces the total energy of the spins. Therefore, the BF, WBF, and GDBF decoding algorithms can be considered deterministic algorithms that determine the most suitable spins to be flipped to reduce the energy $H_{k}^{(X)}(\boldsymbol{x})$ based on the inverse function $\varDelta_{k}^{(X)}(\boldsymbol{x})$. In addition, in contrast to the BF algorithm, the BP algorithm essentially considers soft information, as it calculates a consistent marginal probability $P\left(x_{i}|\boldsymbol{y}\right)$ by exchanging real-valued messages between VNs and CNs. On the other hand, the inversion function and the associated Hamiltonian for the MCMC decoding are formally written by
$\varDelta_{k}^{\mathrm{(MCMC)}}(\boldsymbol{x})=\beta J_{k}x_{k}+\frac{\gamma}{2}\sum_{i\in M(k)}s_{i}(\boldsymbol{x})$ (60)
@@
$H^{\mathrm{(MCMC)}}(\boldsymbol{x})=-\beta\sum_{i=1}^{N_{v}}J_{i}x_{i}+\gamma\sum_{i=1}^{N_{c}}\frac{1-s_{i}(\boldsymbol{x})}{2}.$ (61)
@@
The MCMC decoding algorithm uses MCMC sampling to find $\boldsymbol{x}$ that reduces the energy $H^{\mathrm{(MCMC)}}(\boldsymbol{x})$ based on the inversion function $\varDelta_{k}^{\mathrm{(MCMC)}}(\boldsymbol{x})$. Therefore, the MCMC decoding at the first stage inherently incorporates soft information into the correlation term, thereby resolving the limitation of BF decoding.
@@
Figure 11: Schematic diagrams illustrating two operation modes of the MCMC sampling that optimize (a) MCMC decoding, and (b) MCMC-BF hybrid decoding. The large and small ellipses indicate the search space and the code-state space, respectively. Blue circles indicate the sampled state by the MCMC sampler. Green circles indicate the decoded states by the BF decoding. The red circle indicates the correct code-state $\boldsymbol{z}$. Black solid arrows indicate a sequence of MCMC sampling. Blue broken arrows indicate the map due to the BF decoding.
@@
In addition, MCMC decoding in the first stage also solves another limitation of the BF algorithm in the second stage. The MCMC sampling is stochastic in contrast to our deterministic BF algorithm and gradient descent algorithm used in the GDBF algorithm. This stochasticity introduces randomness into the spin-flip selection. It allows spins to flip even when $\varDelta_{k}^{\mathrm{(MCMC)}}(\boldsymbol{x})>0$, which provides an escape from spurious local minima and makes it more likely to arrive at the neighborhood of the global minimum of $H^{\mathrm{(MCMC)}}(\boldsymbol{x})$. In this way, the stochastic spin-flip selection helps find the global minimum of $H^{\mathrm{(MCMC)}}(\boldsymbol{x})$. A similar stochasticity can be incorporated directly into the GDBF algorithm, either by adding a noise term to the inversion function $\varDelta_{k}^{\mathrm{(GDBF)}}(\boldsymbol{x})$ (Noisy GDBF [29]) or by taking stochastic spin-flip selection into account in the algorithm (Probabilistic GDBF [38]). Briefly speaking, our MCMC-BF hybrid decoding algorithm can be considered an alternative to the BP algorithm and other existing BF decoding algorithms. Our simulation suggests that it is important to properly control the fluctuations of $\boldsymbol{x}$ during MCMC sampling by adjusting the two Lagrange weights $\left\{\beta,\gamma\right\}$ to optimize MCMC-BF decoding. It also suggests that the contribution of the penalty term to the Hamiltonian must be small enough not to force $\boldsymbol{x}$ into a valid code-state. Controlling Lagrange weights is important for switching between the two decoding modes. Fig.11 shows schematic diagrams illustrating two decoding modes: (a) MCMC decoding and (b) MCMC-BF hybrid decoding, respectively. Decoding is formulated as a constrained COP. The large and small ellipses indicate the search space and the code-state space that minimizes the penalty term in the Hamiltonian. In MCMC decoding, the MCMC samples both the code-states (feasible solution) and leakage states (infeasible solution) (Fig.11(a)). In some occasions, the correct code-state (solution) may be sampled. In MCMC-BF hybrid decoding, on the other hand, the first-stage MCMC sampler samples only the leakage state, and the second-stage BF decoding maps the leakage state to the correct code-state occasionally (Fig.11(b)). Our simulations showed that the MCMC-BF hybrid decoding (Fig.11(b)) was over two orders of magnitude more efficient than the MCMC decoding (Fig.11(a)) if we ignore the decoding cost of the post-readout BF decoding and compare performance simply by the number of required MCMC samples needed to obtain at least one correct code-state.
@@ V.2 Spin-flipping mechanism and decoding cost
We briefly discuss the mechanism for selecting which spins to flip at each iteration and the decoding cost of the BF algorithm. Spin-flipping mechanisms are free to choose from a range of strategies. For example, we can flip only the most suitable spin chosen based on the inversion functions $\varDelta_{k}^{(X)}(\boldsymbol{x})$’s estimated by the current decision $\boldsymbol{x}$. Alternatively, we can flip several spins chosen based on $\varDelta_{k}^{(X)}(\boldsymbol{x})$’s at once. These are called the single-spin-flipping and multi-spin-flipping strategies, respectively. In this classification, MCMC decoding algorithm belongs to the single-spin flipping strategy, while our BF decoding algorithm belongs to the multi-spin flipping strategy. Note that the BF decoding algorithm consists of two operations: evaluating the inversion function $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ and determining spins to be flipped. The inversion function $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ for every spin $k$ is computed at once from the current $\boldsymbol{x}$ by matrix multiplication. Subsequent sign evaluation for each element of $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ determines which spins are flipped at once
@@
We can see from Fig.2 that more edges are connected to VNs for the weight-3 syndromes ($d_{v}=K-2$) than the weight-4 syndromes ($d_{v}\leq 4$). Thus, each spin variable affects more syndromes when using the weight-3 syndrome than the weight-4 syndrome. Conversely, each spin variable is determined from more syndromes when using the weight-3 syndrome than the weight-4 syndrome. Consequently, when using the weight-3 syndrome, one must compute the sum of products of data distributed globally across many spins to decide whether to invert each spin. In contrast, when using the weight-4 syndrome, it is sufficient to compute the sum of products of data distributed over at most eight adjacent spins to determine whether to invert each spin. The reader may fear that a global reduction operation to compute the inversion function when using weight-3 syndrome requires more effort at the expense of accelerating error correction at each spin-flip. But this is not the case as discussed below.
@@
We can perform the matrix multiplication required for calculating $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ easily and gain significant benefits from parallel computing techniques and hardware engines that compute matrix sum products, such as GPUs and vector processing engines developed for machine learning. This fact provides practical merit of choosing weight-3 syndrome in $\varDelta_{k}^{\mathrm{(BF)}}(\boldsymbol{x})$ (Eq.(52)). In fact, comparing the computation times of the MCMC and MCMC-BF hybrid decoding presented in Fig.9 performed on the Mathematica platform, the computation time of MCMC-BF hybrid decoding was one-fifth of that of the MCMC decoding. Although the computational cost depends on the algorithm used and the software or hardware platforms on which we perform the calculations, this result suggests that the MCMC-BF hybrid decoding is more efficient than MCMC decoding alone. Furthermore, we confirmed that our BF algorithm was much less computationally expensive than the BP algorithm. For example, the computation time of the performance evaluation for the BF decoding shown in Fig.6 (a) was about 1/40 of that for the BP decoding shown in Fig.6 (b).
@@
This paper does not discuss decoding cost in any further detail. Let us recall that the purpose of this paper is to propose the BF decoding and demonstrate its potential as a practical post-processing algorithm for a pre-processing stochastic algorithm, such as a QA. Since the development of QA devices is ongoing, it is not easy to demonstrate our BF decoding algorithm as a post-processing step for an actual QA device. Instead, the potential of the BF decoding was demonstrated using a classical MCMC sampler as a pre-processor in this study. We believe that the present result depends primarily on the high compatibility between the SLHZ model and the BF decoding algorithm, rather than on the MCMC sampler. Furthermore, our result is consistent with the general belief that two decoding algorithms based on different mechanisms can be combined to solve problems that neither can solve alone. We believe that our BF decoding algorithm is promising for correcting readout errors of the QA devices due to measurement errors as well as dynamic errors that may be encountered during QA.

### v4

@@ V.1 Relationship between MCMC-BF hybrid algorithm and other known BF algorithms
Both the MCMC and BF decoding algorithms are hard-decision algorithms, in which only spin variables are of concern. The MCMC-BF decoding algorithm is considered an extended BF algorithm that incorporates MCMC decoding as a preprocessing step. We discuss the limitations of BF decoding and how they have been resolved in the existing extended BF algorithms.
@@
Many researchers have developed sophisticated versions of the weighted BF (WBF) to incorporate reliability information of the hard decisions into the BF algorithm (see details in Refs.[35, 36] and references therein). Alternatively, there is another approach to incorporating soft information. For example, Wadayama et al. have proposed the gradient-descent bit-flipping (GDBF) algorithm [27]. They are based on modifications to the inversion function that may improve decoding performance, though this comes at the cost of increased decoding complexity. For example, the inversion functions employed in the BF, WBF, and GDBF algorithms are formally written as
$\varDelta_{k}^{\mathrm{(BF)}}(\textbf{x})=1+\sum_{i\in M(k)}s_{i}(\textbf{x}),$ (52)
$\varDelta_{k}^{\mathrm{(WBF)}}(\textbf{x})=\beta|J_{k}|+\sum_{i\in M(k)}w_{k}s_{i}(\textbf{x}),$ (53)
@@
$\varDelta_{k}^{\mathrm{(GDBF)}}(\textbf{x})=J_{k}x_{k}+\sum_{i\in M(k)}s_{i}(\textbf{x}),$ (54)
@@
respectively. Here, $\textbf{x}=(x_{1},\ldots,x_{N_{v}})\in\{\pm 1\}^{N_{v}}$ is the current decision for the spin variables (VNs), $\beta$ is a positive real parameter to be adjusted,
$s_{i}(\textbf{x})=\prod_{j\in N(i)}x_{j}\in\{\pm 1\}$ (55)
@@
is an $i$th syndrome (CN) for the decision x in the spin representation, and $J_{k}$ is identified with channel observation $y_{k}$ in the AWGN channel model. Here, the set $N(i)=\left\{j:H_{ij}=1\right\}$ are the VNs adjacent to an $i$th CN $\left(1\leq i\leq N_{c}\right)$ and the set $M(j)=\left\{i:H_{ij}=1\right\}$ is the CNs adjacent to a $j$th VN $\left(1\leq j\leq N_{v}\right)$. For the definition of these set, see Fig.2. Note that as long as $s_{i}\left(\textbf{x}\right)$ is a weight-3 syndrome, $\varDelta_{k}^{\mathrm{(BF)}}(\textbf{x})$ is symmetric for the permutation of the elements of x. In other words, $\varDelta_{k}^{\mathrm{(BF)}}(\textbf{x})$ is invariant under exchanging $i\longleftrightarrow j$ for any pair of indices of VN other than $k$, justifying the all-one code-state assumption. In contrast, $\varDelta_{k}^{\mathrm{(BF)}}(\textbf{x})$ is not symmetric if $s_{i}\left(\textbf{x}\right)$ is a weight-4 syndrome. Similarly, because $\varDelta_{k}^{\mathrm{(WBF)}}(\textbf{x})$ involves weight $w_{k}$ in the second term and $\varDelta_{k}^{\mathrm{(GDBF)}}(\textbf{x})$ involves $J_{k}$ in the first term, both depend on J. Thus, $\varDelta_{k}^{\mathrm{(WBF)}}(\textbf{x})$ and $\varDelta_{k}^{\mathrm{(GDBF)}}(\textbf{x})$ are not symmetric even if $s_{i}(\textbf{x})$ is a weight-3 syndrome. It is interesting to note that the inversion functions in Eq.(52)-(54) can be formally derived from the following Hamiltonians:
$H^{\mathrm{(BF)}}(\textbf{x})=-\sum_{i=1}^{N_{v}}x_{i}-\sum_{i=1}^{N_{c}}s_{i}(\textbf{x}),$ (56)
$H^{\mathrm{(WBF)}}(\textbf{x})=-\beta\sum_{i=1}^{N_{v}}\left|J_{i}\right|x_{i}-\sum_{i=1}^{N_{c}}w_{i}s_{i}(\textbf{x}),$ (57)
$H^{\mathrm{(GDBF)}}(\textbf{x})=-\frac{1}{2}\sum_{i=1}^{N_{v}}J_{i}x_{i}-\sum_{i=1}^{N_{c}}s_{i}(\textbf{x}),$ (58)
@@
respectively. If we note that when we invert the sign of $x_{k}$, i.e., $x_{k}\rightarrow-x_{k}$, the increase in energy $\Delta H_{k}^{\mathrm{(X)}}(\textbf{x})$ is given by
$\Delta H_{k}^{(X)}(\textbf{x})=2\varDelta_{k}^{(X)}(\textbf{x}),$ (59)
@@
where $X=\mathrm{BF}$, $\mathrm{WBF}$, or $\mathrm{GDBF}$, it follows that if $\varDelta_{k}^{(X)}(\textbf{x})<0$, flipping the $k$th spin reduces the total energy of the spins. Therefore, the BF, WBF, and GDBF decoding algorithms can be considered deterministic algorithms that determine the most suitable spins to be flipped to reduce the energy $H_{k}^{(X)}(\textbf{x})$ based on the inverse function $\varDelta_{k}^{(X)}(\textbf{x})$. In addition, in contrast to the BF algorithm, the BP algorithm essentially considers soft information, as it calculates a consistent marginal probability $P\left(x_{i}|\textbf{y}\right)$ by exchanging real-valued messages between VNs and CNs. On the other hand, the inversion function and the associated Hamiltonian for the MCMC decoding are formally written as
$\varDelta_{k}^{\mathrm{(MCMC)}}(\textbf{x})=\beta J_{k}x_{k}+\frac{\gamma}{2}\sum_{i\in M(k)}s_{i}(\textbf{x})$ (60)
@@
$H^{\mathrm{(MCMC)}}(\textbf{x})=-\beta\sum_{i=1}^{N_{v}}J_{i}x_{i}+\gamma\sum_{i=1}^{N_{c}}\frac{1-s_{i}(\textbf{x})}{2}.$ (61)
@@
The MCMC decoding algorithm uses MCMC sampling to find x that reduces the energy $H^{\mathrm{(MCMC)}}(\textbf{x})$ based on the inversion function $\varDelta_{k}^{\mathrm{(MCMC)}}(\textbf{x})$. Therefore, the MCMC decoding at the first stage inherently incorporates soft information into the correlation term, thereby resolving the limitation of BF decoding.
@@
Figure 11: Schematic diagrams illustrating two operation modes of the MCMC sampling that optimize (a) MCMC decoding, and (b) MCMC-BF hybrid decoding. The large and small ellipses indicate the search space and the code-state space, respectively. The blue circles indicate the sampled state by the MCMC sampler. The green circles indicate the decoded states by the BF decoding. The red circle indicates the correct code state z. The black solid arrows indicate a sequence of MCMC sampling. The blue broken arrows indicate the map due to the BF decoding.
@@
In addition, MCMC decoding in the first stage also solves another limitation of the BF algorithm in the second stage. The MCMC sampling is stochastic in contrast to our deterministic BF algorithm and the gradient-descent algorithm used in the GDBF algorithm. This stochasticity introduces randomness into the spin-flip selection. It allows spins to flip even when $\varDelta_{k}^{\mathrm{(MCMC)}}(\textbf{x})>0$, which provides an escape from spurious local minima and makes it more likely to arrive at the neighborhood of the global minimum of $H^{\mathrm{(MCMC)}}(\textbf{x})$. In this way, the stochastic spin-flip selection helps find the global minimum of $H^{\mathrm{(MCMC)}}(\textbf{x})$. A similar stochasticity can be incorporated directly into the GDBF algorithm, either by adding a noise term to the inversion function $\varDelta_{k}^{\mathrm{(GDBF)}}(\textbf{x})$ (noisy GDBF [28]) or by taking stochastic spin-flip selection into account in the algorithm (probabilistic GDBF [37]). Briefly speaking, our MCMC-BF hybrid decoding algorithm can be considered an alternative to the BP algorithm and other existing BF decoding algorithms. Our simulation suggests that it is important to properly control the fluctuations of x during MCMC sampling by adjusting the two Lagrange weights $\left\{\beta,\gamma\right\}$ to optimize the MCMC-BF decoding. It also suggests that the contribution of the penalty term to the Hamiltonian must be small enough not to force x into a valid code state. Controlling Lagrange weights is important for switching between the two decoding modes. In Fig.11, we show schematic diagrams illustrating two decoding modes: (a) MCMC decoding and (b) MCMC-BF hybrid decoding, respectively. Decoding is formulated as a constrained COP. The large and small ellipses indicate the search space and the code-state space that minimizes the penalty term in the Hamiltonian. In MCMC decoding, the MCMC samples both the code states (feasible solution) and leakage states (infeasible solution) (Fig.11(a)). On some occasions, the correct code state (solution) may be sampled. In MCMC-BF hybrid decoding, on the other hand, the first-stage MCMC sampler samples only the leakage state, and the second-stage BF decoding maps the leakage state to the correct code state occasionally (Fig.11(b)). Our simulations have shown that the MCMC-BF hybrid decoding (Fig.11(b)) is over 2 orders of magnitude more efficient than the MCMC decoding (Fig.11(a)) if we ignore the decoding cost of the postreadout BF decoding and compare performance simply by the number of required MCMC samples needed to obtain at least one correct code state.
@@ V.2 Spin-flipping mechanism and decoding cost
We briefly discuss the mechanism for selecting which spins to flip at each iteration and the decoding cost of the BF algorithm. Spin-flipping mechanisms are free to choose from a range of strategies. For example, we can flip only the most suitable spin chosen based on the inversion function $\varDelta_{k}^{(X)}(\textbf{x})$ values estimated by the current decision x. Alternatively, we can flip several spins chosen based on the $\varDelta_{k}^{(X)}(\textbf{x})$ values at once. These are called the single-spin-flipping and multi-spin-flipping strategies, respectively. In this classification, the MCMC decoding algorithm belongs to the single-spin flipping strategy, while our BF decoding algorithm belongs to the multi-spin-flipping strategy. Note that the BF decoding algorithm consists of two operations: evaluating the inversion function $\varDelta_{k}^{\mathrm{(BF)}}(\textbf{x})$ and determining the spins to be flipped. The inversion function $\varDelta_{k}^{\mathrm{(BF)}}(\textbf{x})$ for every spin $k$ is computed at once from the current x by matrix multiplication. Subsequent sign evaluation for each element of $\varDelta_{k}^{\mathrm{(BF)}}(\textbf{x})$ determines which spins are flipped at once.
@@
We can see from Fig.2 that more edges are connected to VNs for the weight-3 syndromes ($d_{v}=K-2$) than for the weight-4 syndromes ($d_{v}\leq 4$). Thus, each spin variable affects more syndromes when using the weight-3 syndrome than the weight-4 syndrome. Conversely, each spin variable is determined from more syndromes when using the weight-3 syndrome than the weight-4 syndrome. Consequently, when using the weight-3 syndrome, one must compute the sum of products of data distributed globally across many spins to decide whether to invert each spin. In contrast, when using the weight-4 syndrome, it is sufficient to compute the sum of products of data distributed over at most eight adjacent spins to determine whether to invert each spin. The reader may fear that a global reduction operation to compute the inversion function when using weight-3 syndrome will require more effort at the expense of accelerating error correction at each spin flip, but this is not the case, as discussed below.
@@
We can perform the matrix multiplication required for calculating $\varDelta_{k}^{\mathrm{(BF)}}(\textbf{x})$ easily and gain significant benefits from parallel-computing techniques and hardware engines that compute matrix sum products, such as GPUs and vector processing engines developed for machine learning. This fact provides the practical merit of choosing the weight-3 syndrome in $\varDelta_{k}^{\mathrm{(BF)}}(\textbf{x})$ (Eq.(52)). In fact, comparing the computation times of the MCMC and MCMC-BF hybrid decoding presented in Fig.9 performed on the Mathematica® platform, the computation time of MCMC-BF hybrid decoding was one fifth of that of the MCMC decoding. Although the computational cost depends on the algorithm used and the software or hardware platforms on which we perform the calculations, this result suggests that the MCMC-BF hybrid decoding is more efficient than MCMC decoding alone. Furthermore, we confirmed that our BF algorithm was much less computationally expensive than the BP algorithm. For example, the computation time of the performance evaluation for the BF decoding shown in Fig.6 (a) was about 1/40 of that for the BP decoding shown in Fig.6 (b).
@@
This paper does not discuss decoding cost in any further detail. Let us recall that the purpose of this paper is to propose the BF decoding and demonstrate its potential as a practical postprocessing algorithm for a preprocessing stochastic algorithm, such as a QA. Since the development of QA devices is ongoing, it is not easy to demonstrate our BF decoding algorithm as a postprocessing step for an actual QA device. Instead, the potential of the BF decoding has been demonstrated using a classical MCMC sampler as a preprocessor in this study. We believe that the present result depends primarily on the high compatibility between the SLHZ model and the BF decoding algorithm, rather than on the MCMC sampler. Furthermore, our result is consistent with the general belief that two decoding algorithms based on different mechanisms can be combined to solve problems that neither can solve alone. We believe that our BF decoding algorithm is promising for correcting readout errors of the QA devices due to measurement errors as well as dynamic errors that may be encountered during QA.